# Restyling Unsupervised Concept Based Interpretable Networks with Generative Models

### v1

@@ 0.B.2.3 Details for VisCoIN
0.B.3 Evaluation details
@@
Metric computation . The median faithfulness is computed over 1000 random samples from the test data. For consistency, we use $N_{cc}=100,\lambda=2$, i.e., given any concept $\phi_{k}$, we extract its 100 most activating samples over samples of classes its most relevant for. The constant high activation is twice ($\lambda=2$) the maximum activation of $\phi_{k}(x)$ over the pool of 100 samples. Thus the binary training dataset created via samples from training data consists of 200 samples, 100 “positive” samples with high activation of $\phi_{k}(x)$ and 100 “negative” samples with zero activation $\phi_{k}(x)$. The binary testing dataset also contains the same number of samples of each type but is created via samples from test data. The feature maps we extract are the output of the second block of the pretrained $f$ (ResNet50), the 22nd convolutional layer. The shape for each feature map is $512\times 32\times 32$. We pool them across the spatial axis to obtain an embedding of size $512$ for any input sample. The linear classifier we train is a linear SVM. We select its inverse regularization strength $C$ from the set $\{0.01,0.1,1.0,5.0\}$ (lower value is stronger regularization). The parameter is selected using 5-fold cross validation on the created training data.
0.B.3.1 Baseline implementations
@@
We utilize the official codebase available for FLINThttps://github.com/jayneelparekh/FLINT and FLAEMhttps://github.com/anirbansarkar-cs/Ante-hoc_Explainability_Concepts for our baseline implementation. For fairness, we use the same number of concepts for both of them. Since our architecture is closer to FLINT, we update and adapt it to implement in similar settings as ours. We use the same $f$ architecture (ResNet50 instead of ResNet18) for both the systems and keep it pretrained and fixed. The $\Psi$ architecture is also similar in that it accesses the same set of hidden layers and has the same structure and depth. For other hyperparameters we use their default settings applied earlier for CUB-200. Implementing FLAEM with same network architecture is more complicated as it deviates considerably from the proposed architecture, thus we mostly use their default settings. In their code, they use a base classifier architecture similar to ResNet101 and use the output of final conv layer as the concept representation. In both cases, we do not modify the decoders. FLAEM uses a simpler decoder that learns 3 deconvolution layers, while FLINT learns a deeper decoder consisting of transposed convolution layers.
Appendix 0.C Additional experiments
0.C.1 Consistency with higher $\lambda$
@@
We present the results here for evaluating consistency with a higher value of $\lambda=3$, compared to the main text where $\lambda=2$. Thus, the constant high activation of $\phi_{k}(x)$ used to generate “positive” samples of the dataset is increased further. We thus expect the “separation” in the embedding space to increase and consequently a higher performance $CC_{k}$ of the binary classifier $\varphi_{k}$ for any $k$. The results for both are presented in Tab. 4. The results confirm that emphasizing the concept indeed makes the visual modifications more stronger and consistent. Moreover, we also observe that increase in consistency of VisCoIN tends to be larger than increase for other CoIN systems.
0.C.2 Qualitative visualization FID
Table 5: Quantitative evaluation (FID) of the visualization obtained for interpretation. Lower is better.
@@ 0.C.2 Qualitative visualization FID
While qualitatively on can clearly observe the difficulty to understand activation maximization based visualization in FLINT. We further support our claim about the unnaturalness of these visualizations compared to visualization in VisCoIN by computing the distance of distributions of visualizations in FLINT, visualizations in VisCoIN and original data distribution. Note that we can’t use any reconstruction metrics as FLINT visualizations don’t reconstruct a given input. Instead they initialize using a given maximum activating sample and execute the optimization procedure of activation maximization to maximally activate a $\phi_{k}(x)$. We thus compute the FID distance between the visualizations and the data distribution. For VisCoIN visualization we select our most extreme value of $\lambda$. For FLINT visualization we follow their implementation and run the input optimization procedure for 1000 iterations. Since the FLINT visualizations are relatively lot more expensive to compute (1000 backward passes vs 1 forward pass for VisCoIN), we compute the visualizations for 3 maximum activating samples for random 400 relevant class-concept pairs (with $r_{k,c}>0.5$). Thus the FIDs are computed between 1200 data samples and corresponding visualizations.
Appendix 0.D Ablation studies
@@
We present ablation studies for components and simultaneously discuss our rationale behind the design selection of these components, (a) effect of reconstruction-classification loss with weight $\gamma$, (b) role of using a supporting representation $\Phi^{\prime}(x)$ to assist in reconstruction, (c) selection of number of concepts $K$, (d) effect of orthogonality loss, and (e) effect of fidelity and sparsity loss weights.
@@
We highlight at this point to the reader that there is an overarching theme in our various design choices. Most of them are based on shaping the systems suitability for better optimization of perceptual similarity for reconstruction. We constantly aim to achieve better reconstruction without major negative impacts for any other properties. This is because for complex datasets (CUB-200, Stanford Cars), the key bottleneck in the design is to achieve the high-quality reconstruction for viewability.
@@
Table 6: Effect of the weight $\gamma$ of the Reconstruction-Classification loss in the total training loss, measured by Faithfulness, LPIPS and FID, on CUB-200. Faithfulness computed with a threshold of 0.2. Bold indicates setting selected for our experiments.
$\gamma$  Faithfulness  MSE  LPIPS  FID
@@
$\Phi^{\prime}$  K  MSE  LPIPS  FID
@@
0.D.0.1 Reconstruction-classification loss
@@
Tab. 6 reports the perceptual similarity and faithfulness with threshold $\tau=0.2$ on test data of CUB-200 for different strength $\gamma$. Interestingly, it indicates a tradeoff between faithfulness and perceptual similarity. Completely removing this loss heavily impacts the faithfulness. However, among the positive $\gamma$, our choice was driven strongly by achieving a reconstruction with high-enough quality and perceptual similarity to enable effective visualization. Thus, we chose $\gamma=0.1$. The key reason for this is that perceptual similarity is a much better indicator for viewability. For instance, for the $\gamma=0.2$, even though the FID is better, the reconstruction (LPIPS) is noticeably worse. Reconstruction of the final training batch is indicated in 5, to highlight this issue with high $\gamma$. We thus kept a smaller $\gamma$ for CUB-200 and Stanford Cars where the reconstruction is more challenging and slightly higher value for CelebA-HQ.
0.D.0.2 Use of support representation $\Phi^{\prime}$
Tab. 7 presents the reconstruction metrics for different concept dictionary sizes and use of $\Phi^{\prime}$ in reconstruction. As before, we prioritized optimization of LPIPS and using $\Phi^{\prime}$ assists in achieving better reconstruction whilst allowing us to employ a smaller dictionary.
0.D.0.3 Selecting number of concepts $K$
The ablation with different number of concepts is given in Tab. 8 and is an important hyperparameter of the system. While choosing $K$, it is easy to filter out smaller $K$ values as they clearly lead to a worse reconstruction. However, hypothetically a higher $K$ should improve for all the metrics and thus finding an upper bound for $K$ is more subjective. Our choice was mainly influenced by (1) the observation that increasing from 256 to 512 offered relatively minimal advantage in reconstruction, and (2) previous methods that used supervised concepts train with 312 concepts. Hence we intended to use a similar dictionary size. Since the Stanford Cars dataset had a comparable number of samples and classes, we used the same number of concepts. For CelebA-HQ, we experimented with reduced $K$ as there are only two classes and the images are less diverse compared to the other two datasets.

### v2

@@ C.2.3 Training VisCoIN
C.3 Evaluation details
C.3.1 Metric computation
@@
The median faithfulness is computed over 1000 random samples from the test data. For consistency, we use $N_{cc}=100,\lambda=2$, i.e., given any concept $\phi_{k}$, we extract its 100 most activating samples over samples of classes its most relevant for. The constant high activation is twice ($\lambda=2$) the maximum activation of $\phi_{k}(x)$ over the pool of 100 samples. Thus the binary training dataset created via samples from training data consists of 200 samples, 100 “positive” samples with high activation of $\phi_{k}(x)$ and 100 “negative” samples with zero activation $\phi_{k}(x)$. The binary testing dataset also contains the same number of samples of each type but is created via samples from test data. The feature maps we extract are the output of the second block of the pretrained $f$ (ResNet50), the 22nd convolutional layer. The shape for each feature map is $512\times 32\times 32$. We pool them across the spatial axis to obtain an embedding of size $512$ for any input sample. The linear classifier we train is a linear SVM. We select its inverse regularization strength $C$ from the set $\{0.01,0.1,1.0,5.0\}$ (lower value is stronger regularization). The parameter is selected using 5-fold cross validation on the created training data.
C.3.2 Baseline implementations
@@
We utilize the official codebase available for FLINThttps://github.com/jayneelparekh/FLINT and FLAEMhttps://github.com/anirbansarkar-cs/Ante-hoc_Explainability_Concepts for our baseline implementation. For fairness, we use the same number of concepts for both of them. Since our architecture is closer to FLINT, we update and adapt it to implement in similar settings as ours. We use the same $f$ architecture for both the systems and keep it pretrained and fixed. The $\Psi$ architecture is also similar in that it accesses the same set of hidden layers and has the same structure and depth. For other hyperparameters we use their default settings applied earlier for CUB-200. Implementing FLAEM with same network architecture is more complicated as it deviates considerably from the proposed architecture, thus we mostly use their default settings. In their code, they use a base classifier architecture similar to ResNet101 and use the output of final conv layer as the concept representation. In both cases, we do not modify the decoders. FLAEM uses a simpler decoder that learns 3 deconvolution layers, while FLINT learns a deeper decoder consisting of transposed convolution layers.
Appendix D Experiments with other architectures for $f$ and $G$
In this section, we demonstrate that our model can generalize to a variety of network architectures for $f$ and $G$. In particular, we show applicability of VisCoIN with LeNet, ResNet101 and ViT-B/16 architectures for $f$, and $\beta$-VAE and ProgressiveGANs architectures for $G$.
[…]
Method Acc. ($\uparrow$) MSE ($\downarrow$) LPIPS ($\downarrow$) FID ($\downarrow$) $FF_{x}$ ($\uparrow$) $CC_{k}$ ($\uparrow$)
VisCoIN - ProgressiveGAN 87.82 0.095 0.453 6.98 0.37 93.8 $\pm$ 6.8
VisCoIN - StyleGAN2-ADA 87.71 0.094 0.405 8.55 0.171 85.5 $\pm$ 13.9
Figure 5: Qualitative results of VisCoIN with ProgressiveGAN (Karras et al., 2018) on CelebA-HQ dataset.
[…]
Method Acc. ($\uparrow$) MSE ($\downarrow$) $FF_{x}$ ($\uparrow$) $CC_{k}$ ($\uparrow$)
0.4 0.2 0.1
VisCoIN - $\beta$-VAE 88.06 0.029 0.576 0.693 0.762 94.3 $\pm$ 7.4
FLINT 86.41 0.031 0.564 0.688 0.728 96.2 $\pm$ 4.8
D.1 Experiment with ProgressiveGAN
We present in Table 4, results of experiments using a pretrained ProgressiveGAN (Karras et al., 2018) as our generative model $G$, on CelebA-HQ dataset, all other hyperparameters being identical as experiments presented in the main text on this dataset. As can be seen in the table, we obtain similar accuracy and MSE, but FID, faithfulness and consistency are better using ProgressiveGAN, while LPIPS is better using StyleGAN2-ADA. We show an example concept visualization in Fig. 5. The visualization clearly reveal two features (‘Smile’, ‘Long Hair’) both controlled by a single concept function. The high consistency and faithfulness is also possibly the result of strong modifications in generated images when traversing latent space in ProgressiveGAN. However, most importantly, reconstruction quality worsens compared to results on StyleGAN and affects visualization. This is a major factor in our preference of the StyleGAN version of VisCoIN over ProgressiveGAN version of VisCoIN. This is also a good example to illustrate the important role viewability plays in visualization.
D.2 Experiment with Beta-VAE
We also experiment with applying VisCoIN using $\beta$-VAE (Higgins et al., 2017) as generative model, on FashionMNIST dataset (Xiao et al., 2017). We use $K=25$ as the number of concepts. The VAE is trained with $\beta=2$ and latent embedding size of 48. Given that the dataset consists of grayscale images of smaller resolution, we do not use LPIPS loss during training. We use a $f$ architecture similar to LeNet (LeCun, 2015).
Table 5 reports the result of our VisCoIN system against FLINT as a baseline. For reconstruction, we only report the MSE since both LPIPS and FID rely on fixed networks pretrained on large scale color images and are thus not suitable for this dataset. We can see that our VisCoIN achieves better accuracy, reconstruction quality and faithfulness than FLINT. For small and grayscale images, even though we observe an advantage in using a generative model over a standard decoder, the effects are less pronounced compared to more complex images and classification tasks (as in main text). The main reason for this is that in the current scenario, the standard decoders are also quite capable at generating viewable reconstructions. This gap with generative models grows larger as the underlying data distributions grows more complex. Nonetheless, crucially, these experiments show that our method is effective even with a completely different generative model. Qualitative results of two concepts and their visualizations can be found in Fig. 6.
(a) Concept “Heel” in class Ankle boot
(b) Concept “Cut between legs” in class Trousers
Figure 6: Qualitative results for VisCoIN - $\beta$-VAE on FashionMNIST.
D.3 Experiment with ResNet101
[…]
Method Acc. ($\uparrow$) MSE ($\downarrow$) LPIPS ($\downarrow$) FID ($\downarrow$) $FF_{x}$ ($\uparrow$) $CC_{k}$ ($\uparrow$)
VisCoIN - RN50 79.44 0.16 0.545 15.85 0.146 85.0 $\pm$ 8.4
VisCoIN - RN101 79.44 0.16 0.542 11.24 0.202 94.0 $\pm$ 6.2
We present in Table 6 additional experiments on CUB dataset, when using a ResNet101 network as our classification network $f$, all other hyperparameters being identical. We can see that VisCoIN achieves similar accuracy but better perceptual reconstruction quality (LPIPS and FID), faithfulness and consistency. While relatively nominal, the consistent improvement in all interpretability evaluation metrics compared to ResNet-50 is likely the result of better quality of hidden layers supplied by a larger backbone.
D.4 Experiment with ViT-B/16
We report in Table 7 an additional experiment using a ViT-B/16 for $f$ on CUB dataset, while keeping other architectures almost identical to experiments in main text. We started from a ViT-B/16, pretrained on ImageNet-21K (Wu et al., 2020; Deng et al., 2009) and only finetuned the classification head on CUB, to use as $f$. We take the patch embeddings of final layer as input to $\Psi$. We keep identical $\Omega$, $\Theta$ hyperparameters, and slightly modify $\Psi$ for reduced number of feature maps. As can be seen from the numerical results below, we achieve better accuracy thanks to a better pretrained $f$, but reconstruction (LPIPS) and faithfulness are worse. The results could be improved by better designing $\Psi$ and accessing more internals embeddings. However, they certainly show that VisCoIN can generalize to other backbone architectures. An example concept visualization to support this is shown in Fig. 7.
Figure 7: Example concept visualization ($\phi_{36}$, class 15) for ViT-B/16 on CUB-200 (”Red front/belly”). We show the original input, generated/reconstructed images with intervened activations ($\lambda=0,4$) and the difference between generated images to localize the modifications.
[…]
Model Acc. $f$ Acc. $g$ LPIPS ($\downarrow$) FID ($\downarrow$) $FF_{x}$ ($\tau=0.2$) ($\uparrow$)
VisCoIN - ResNet50 80.56 79.44 0.545 15.85 0.146
VisCoIN - ViT-B/16 86.66 85.86 0.582 14.88 0.081
Appendix E Additional evaluations
E.1 Consistency with higher $\lambda$
@@
We present the results here for evaluating consistency with a higher value of $\lambda=3$, compared to the main text where $\lambda=2$. Thus, the constant high activation of $\phi_{k}(x)$ used to generate “positive” samples of the dataset is increased further. We thus expect the “separation” in the embedding space to increase and consequently a higher performance $CC_{k}$ of the binary classifier $\varphi_{k}$ for any $k$. The results for both are presented in Table 8. The results confirm that emphasizing the concept indeed makes the visual modifications more stronger and consistent. Moreover, we also observe that increase in consistency of VisCoIN tends to be larger than increase for other CoIN systems.
E.2 Qualitative visualization FID
Table 9: Quantitative evaluation (FID) of the visualization obtained for interpretation with the original data distribution (1200 samples). Lower is better. For FLINT, visualization is activation maximization output. For VisCoIN, visualization is generated image with $\lambda=4$.
@@ E.2 Qualitative visualization FID
While qualitatively, one can clearly observe the difficulty to understand activation maximization based visualization in FLINT. We further support our claim about the unnaturalness of these visualizations compared to visualization in VisCoIN by computing the distance of distributions of visualizations in FLINT, visualizations in VisCoIN and original data distribution, reported in Table 9. Note that we can’t use any reconstruction metrics as FLINT visualizations don’t reconstruct a given input. Instead they initialize using a given maximum activating sample and execute the optimization procedure of activation maximization to maximally activate a $\phi_{k}(x)$. We thus compute the FID distance between the visualizations and the data distribution. For VisCoIN visualization we select our most extreme value of $\lambda$. For FLINT visualization we follow their implementation and run the input optimization procedure for 1000 iterations. Since the FLINT visualizations are relatively lot more expensive to compute (1000 backward passes vs 1 forward pass for VisCoIN), we compute the visualizations for 3 maximum activating samples for random 400 relevant class-concept pairs (with $r_{k,c}>0.5$). Thus the FIDs are computed between 1200 data samples and corresponding visualizations.
E.3 Top-N concept activation filter
To preserve maximal amount of interpretable by-design structure, we design the $\Theta$ function as a single linear layer with softmax, similar to other unsupervised CoINs. Nevertheless, this by-design interpretable structure also “erodes” as the size of concept dictionary $K$ increases. For large dictionary sizes, if the activations are not extremely sparse, interpreting the prediction even through a linear $\Theta$ can become tedious and less interpretable (Lipton, 2018).
One way to preserve the interpretable structure even with large $K$ is by controlling the number of non-zero concept activations. We experimented applying a “Top-$N$” function on $\Phi(x)$ before $\Theta$, to keep only the most activated concept for prediction, for different values of $N$, and report results in Table 10. Although it improves interpretability and conciseness of interesting concepts, it comes at the cost of accuracy of the overall system. However, we can see that using about 25% of the most activated concepts still preserves good accuracy in general.
[…]
Dataset $N$
$4$ $8$ $16$ $32$ $64$ $128$ $256$
CUB 23.75 42.25 59.28 70.07 76.25 78.97 79.44
Stanford Cars 13.38 26.43 45.75 62.76 72.43 77.20 79.89
CelebA-HQ 79.92 80.63 84.00 86.90 87.71
E.4 AUC-Faithfulness metric
We performed preliminary experiments to compare faithfulness of VisCoIN and FLINT on CUB-200 by computing the area under the curve (AUC) of accuracy w.r.t number of “added” concepts.
Specifically, for each test sample $x$, we initialize $\Phi(x)$ to 00, increasingly add the most activated top-$N$ concepts for $N\in\{2,4,8,16,32,64,128,200,256\}$ to $\Phi(x)$, plot the accuracy of $g(x^{\prime}),x^{\prime}=G(\Omega(\Phi(x)_{N}))$ and compute its AUC.
We report the AUC in Table 11. Note that since the accuracy is on generated images, the accuracy of $g$ is lower than its accuracy on the same dataset. The results are strongly in favor of VisCoIN. We expect VisCoIN to generally outperform other unsupervised CoINs on this metric as it’s capable to generate high-quality reconstructions.
Table 11: AUC of accuracy curve when gradually adding most activated top-$N$ ($N\in\{2,4,8,16,32,64,128,200,256\}$ concepts for each test sample to $\Phi(x)$, along with final accuracy of $g$ (in %) on reconstructed images, on CUB-200 dataset. Higher is better.
Method AUC-FF metric Acc. $g$ on reconstructions
VisCoIN 0.407 58%
FLINT 0.042 4.5%
E.5 Additional consistency evaluation
For datasets with annotations for presence/absence of different object parts, consistency for any given concept function $\phi_{k}$ can also be evaluated by analyzing the “consistency” of their impact on the “presence” of object parts in the generated images. Among the three tasks we evaluate VisCoIN on, the CUB-200 dataset consists of binary annotations for 312 bird-parts for each image. Since such annotations are not available for generated images, prior to the evaluation, we first train a ResNet-50 based predictor to predict probabilities for each bird part. Let this predictor be denoted as $f_{part}$.
For a given concept function $\phi_{k}$ the goal of this evaluation is to ascertain if there is a bird part which is “consistently affected” when activation $\phi_{k}(x)$ is modified. To quantify this notion, we first select test images of the classes for which $\phi_{k}$ is highly relevant, i.e. $r_{k,c}>0.5$. This set of images is denoted as $\mathcal{X}_{k}$. For each $x\in\mathcal{X}_{k}$, we compute $\Phi(x)$ and generate two images with intervened activations, one with $\phi_{k}(x)=0$ and other with $\phi_{k}(x)=\lambda\times\phi_{k}^{max}$ as done for consistency evaluation in main paper. We use $\lambda=3$ for this experiment. The generated images are denoted as $x_{k}^{-},x_{k}^{+}$ respectively. Note that this dataset preparation is very similar to one used for consistency evaluation in main paper.
We compute part probabilities for all 312 parts as $f_{part}(x_{k}^{-}),f_{part}(x_{k}^{+})$. We define $\phi_{k}$ to significantly affect part $j$ for image $x$ if the probability prediction of part $j$ increases when $\phi_{k}$ activates by more than threshold $\tau$, i.e. $f_{part}(x_{k}^{+})_{j}-f_{part}(x_{k}^{-})_{j}>\tau$. For each bird part $j$, its part consistency is computed as fraction of $x\in\mathcal{X}_{k}$ s.t. $\phi_{k}$ significantly affects part $j$ in $x$. Consistency of $\phi_{k}$ is calculated as maximum part consistency over all the parts
$CC_{k}^{part}(\mathcal{X}_{k};\tau)=\max_{j}\frac{|\{x:f_{part}(x_{k}^{+})_{j}-f_{part}(x_{k}^{-})_{j}>\tau\}|}{|\mathcal{X}_{k}|}$
We report results for mean of $CC_{k}^{part}(\mathcal{X}_{k};\tau)$ with $\tau\in\{0.05,0.1,0.2\}$ for FLINT and VisCoIN in Table 12. We also report the fraction of concepts that significantly affect some bird part for more than 50% of the images in Fig. 8. Both results indicate that on average concepts in VisCoIN affect some bird part, more consistently, compared to FLINT.
Table 12: Average part consistency over all concepts for FLINT, VisCoIN on CUB-200 (Scale: 0-1, Higher is better).
Method Part probability change threshold $\tau$
$\tau=$ 0.05 $\tau=$ 0.1 $\tau=$ 0.2
VisCoIN 77.7 $\pm$ 13.2 64.4 $\pm$ 16.1 41.7 $\pm$ 15.7
FLINT 56.6 $\pm$ 27.5 39.6 $\pm$ 27.1 18.8 $\pm$ 20.8
Figure 8: Fraction of concepts which affect some bird part significantly for more than 50% of their relevant images ($CC_{k}^{part}(\mathcal{X}_{k};\tau)>0.5$). Results reported for different levels of significant part probability change threshold $\tau$ for VisCoIN, FLINT on CUB-200.
E.6 Sparsity evaluation
We report in Table 13 sparsity results of VisCoIN and FLINT for CUB-200 at different relevance thresholds. The sparsity is calculated as the average number of relevant concepts per class such that global relevance $r_{k,c}>\text{threshold}$.
FLINT achieves better sparsity because of its use of entropy based losses to compress $\Phi(x)$. While sparsity of VisCoIN could be improved by increasing the $L_{1}$ regularization weight, we prioritized optimizing for reconstruction/viewability because (i) For previous unsupervised CoINs this is a major limitation, (ii) The current levels of sparsity seemed reasonable (total number of concepts $K=256$ is much higher than relevant for any class), (iii) Excessive compression of information can make concepts less interpretable and similar to class logits.
Table 13: Sparsity, measured as average number of concepts per class with global relevance higher than a threshold, of VisCoIN and FLINT on CUB-200 dataset. Lower is better.
Method Threshold on $r_{k,c}$
 0.7  0.5  0.2
VisCoIN 2.3 6.1 27.1
FLINT 1.5 3.4 10.2
Appendix F Ablation studies
@@
We present ablation studies for various components of our system and simultaneously discuss our rationale behind the design selection of these components. Specifically, we study (a) effect of reconstruction-classification loss with weight $\gamma$ in Section F.1, (b) role of using a supporting representation $\Phi^{\prime}(x)$ to assist in reconstruction in Section F.2, (c) selection of number of concepts $K$ in Section F.3, (d) effect of orthogonality loss in Section F.4, (e) effect of fidelity and sparsity loss weights in Section F.5, and (f) usefulness of concept translator $\Omega$ in Section F.6.
@@
We highlight at this point to the reader that there is an overarching theme that governed many of our design choices. In particular, most of them are based on shaping the systems suitability for better optimization of perceptual similarity for reconstruction. We constantly aim to achieve better reconstruction without major negative impacts for any other properties. This is because for complex datasets (CUB-200, Stanford Cars), the key bottleneck in the design is to achieve the high-quality reconstruction for viewability.
@@
Table 14: Effect of the weight $\gamma$ of the Reconstruction-Classification loss in the total training loss, measured by Faithfulness, LPIPS and FID, on CUB-200. Faithfulness computed with a threshold of 0.2. Bold indicates setting selected for our experiments. Experimentally, $\gamma$ causes tradeoff between faithfulness and quality of reconstruction.
$\gamma$  Faithfulness ($\uparrow$)  MSE ($\downarrow$)  LPIPS ($\downarrow$)  FID ($\downarrow$)
@@
$\Phi^{\prime}$ K  MSE ($\downarrow$)  LPIPS ($\downarrow$)  FID ($\downarrow$)
@@
F.1 Reconstruction-classification loss
@@
Table 14 reports the perceptual similarity and faithfulness with threshold $\tau=0.2$ on test data of CUB-200 for different strength $\gamma$. Interestingly, it indicates a tradeoff between faithfulness and perceptual similarity. One possible reason for this tradeoff is that a higher $\gamma$ can push the model to generate more spurious features captured by the classifier at the expense of input quality. Thus, even though the reconstructed images remain relatively far from input this can still lead to “accurate” predictions from the classifier i.e. it predicts the same class as input. Completely removing this loss heavily impacts the faithfulness. However, among the positive $\gamma$, our choice was driven strongly by achieving a reconstruction with high-enough quality and perceptual similarity to enable effective visualization. Thus, we chose $\gamma=0.1$. The key reason for this is that perceptual similarity is a much better indicator for viewability. For instance, for the $\gamma=0.2$, even though the FID is better, the reconstruction (LPIPS) is noticeably worse. Reconstruction of the final training batch is indicated in 9, to highlight this issue with high $\gamma$. As is apparent from the figure, the perceptual similarity and consequently the viewability of the model is poor with high $\gamma$. We thus kept a smaller $\gamma$ for CUB-200 and Stanford Cars where the reconstruction is more challenging and slightly higher value for CelebA-HQ.
F.2 Use of support representation $\Phi^{\prime}$
The construct of support representation and our use of it is limited to StyleGAN as architecture of $G$. We describe its precise design leveraging the StyleGAN architecture in Section C.1.1. While it is not essential to learning and operation of VisCoIN, its inclusion offers a lever to achieve better reconstruction without having to increase the concept dictionary size $K$. Table 15 presents the reconstruction metrics for different concept dictionary sizes and use of $\Phi^{\prime}$ in reconstruction. Using $K=256$ with the support representation offers a slightly better reconstruction than even using $K=512$ but no support representation. As before, we prioritized optimization of LPIPS and using $\Phi^{\prime}$ assists in achieving better reconstruction whilst allowing us to employ a smaller dictionary.
F.3 Selecting number of concepts $K$
The ablation with different number of concepts is given in Table 16 and is an important hyperparameter of the system. While choosing $K$, it is easy to filter out smaller $K$ values as they clearly lead to a worse reconstruction. However, hypothetically a higher $K$ should improve for all the metrics since it allows for more expressivity in the concept representation. Thus finding an upper bound for $K$ is more subjective. Our choice was mainly influenced by (1) the observation that increasing from 256 to 512 offered relatively minimal advantage in reconstruction, and (2) previous methods that used supervised concepts train with 312 concepts. Hence we intended to use a similar dictionary size. Since the Stanford Cars dataset had a comparable number of samples and classes, we used the same number of concepts. For CelebA-HQ, we experimented with reduced $K$ as there are only two classes and the images are less diverse compared to the other two datasets.